# Understanding and Estimating the Execution Time of Quantum Programs

### v1

@@ 2.3 Classical software execution time estimation
3 Background
3.1 Quantum circuits
Fig. 1 shows a simple quantum circuit that represents a quantum program that we developed using IBM’s open-source software development toolkit, Qiskit. A quantum circuit typically consists of a combination of quantum bits (qubits), classical bits, quantum gates, and measurement gates. In the example shown in Fig. 1, $q_{0}$ and $q_{1}$ represent two qubits, whereas $c$ represent two classical bits. Unlike classical bits, each qubit can exist simultaneously in both $|0\rangle$ and $|1\rangle$ states.
@@ 3.1 Quantum circuits
  • The blue box with “H” symbol: Hadamard gate, applied to $q_{0}$.
  • The blue box with “+” symbol: CNOT gate, with $q_{0}$ as the control qubit and $q_{1}$ as the target qubit.
@@
Figure 1: An example of a Quantum Circuit
3.2 Quantum circuit compilation
@@
Like the classical domain, implementing a quantum algorithm on a real backend involves transforming it into a format that meets the specific hardware constraints [19]. This transformation encompasses four stages: the algorithmic level, target-independent level, target-dependent native gates level, and target-dependent mapped level.
@@
Quantum algorithms are typically designed and trialed at a hardware-neutral stage, known as the algorithmic level [19]. Here, they are expressed as quantum circuits made up of high-level elements, free of restrictions related to specific gate-sets or hardware designs.
@@
The initial step in applying a conceptual quantum algorithm to a specific problem is to create and refine these high-level components, disregarding the intended hardware. This process is akin to the function of a classical compiler and includes optimizing and simplifying the representation, now termed the target-independent level [19].
@@
Modern quantum backends have specific limitations, necessitating a target-dependent compilation phase. Backends offer a limited range of native gates, usually an entangling gate and a set of single-qubit gates. Circuits must therefore be adapted to this native gate-set and optimized to minimize additional complexity. This adaptation is known as the target-dependent native gates level [19].
@@
Additionally, current backends, particularly those using superconducting qubits, have restricted qubit connectivity. This requires mapping the circuit’s logical qubits to the backend’s physical qubits, ensuring that multi-qubit gates apply only to directly connected qubits. This dynamic mapping, followed by optimization to reduce compilation overhead, is defined as the target-dependent mapped level [19].
@@
Finally, the circuit is prepared for execution on the quantum computer. It involves scheduling the gates, calibrating, and converting the circuit into machine code [19]. The final binary code is then executed on the backend to perform the quantum computation.
3.3 Quantum execution and execution time
@@
When a quantum circuit is submitted for execution on a quantum computing platform such as the IBM Quantum Platform, the platform usually provides an estimated execution time. The execution time includes loading the quantum circuit into the pre-selected quantum computer that creates the pulses, running multiple repetitions of calibration pulses and the quantum circuit, and resetting the qubits (relaxation) along with either calibration or executing the quantum circuit. These steps together make up the total execution time. As illustrated in Fig. 2(a), on the IBM quantum computer ibm_osaka, the quantum circuit is expected to run for 2 minutes and 27.5 seconds. After the execution, the platform provides the actual execution time of the circuit. As shown in Fig. 2(b), the actual execution time is only 27 seconds, substantially different from the estimation. The primary objective of our approach is to accurately predict the execution time, aiming to substantially reduce the significant errors observed in the estimates provided by quantum computing platforms. While the IBM Quantum Platform serves as a case study in this paper, our goal extends beyond this to a more general application across various platforms.
@@
Figure 2: Estimated and Actual Execution Times on IBM Quantum Platform
4 Data Preparation
In this section, we describe our data preparation process, including the selection of the quantum programs and quantum backends, as well as the measurement of execution time.
4.1 Quantum program selection
We select our subject quantum programs from the MQT Benchmark [18] which provides an extensive collection of quantum programs covering various quantum computing applications across different levels of abstraction. This library comprises 3,000 quantum programs, spanning from 2 to 130 qubits. Notably, this dataset exhibits significant diversity as it considers two qubit technologies, seven different backends, two compilers, and six corresponding settings. This diversity renders it a robust and versatile resource for our research that aims to estimate quantum execution time for diverse quantum programs. Due to the limitation of quantum computing resources, we must select a subset for our experiments. For the compiler, we have chosen Qiskit, an open-source SDK tailored for quantum computing. Qiskit specializes in circuit, pulse, and algorithm levels, enabling the creation and manipulation of quantum programs. These programs can be executed on prototype quantum backends available through the IBM Quantum Platform or via simulators on local machines. Adhering to the circuit model of universal quantum computation, Qiskit is versatile, supporting various quantum hardware types that comply with this model, including superconducting qubits and trapped ions.
@@ 4.1 Quantum program selection
Finally, as for the quantum programs, we selected all quantum programs ranging from 2 qubits to 127 qubits (the limit of our considered quantum backends). When executing these quantum programs, we found that a small portion encountered loading or compilation errors, and some had execution times that were excessively long, exceeding 10 minutes. Therefore, we excluded these quantum programs with errors and set a timeout function to filter out those with excessively long execution times (over 10 minutes). Ultimately, we obtained a total of 1,510 subject quantum programs.
4.2 Quantum backend selection
We consider the quantum backends provided by the IBM Quantum Platform as it is one of the widely used quantum computing platforms and provides open accesses for free. For open accesses, the accessible quantum computers include ibm_brisbane, ibm_osaka, and ibm_kyoto (at the time of writing), all of which support 127 physical qubits. Due to the high traffic on ibm_brisbane and in order to ensure data diversity, we have chosen to execute the quantum circuits on both ibm_osaka and ibm_kyoto. To maintain consistency in data dimensions while leveraging simulation data and active learning to select samples for execution on real backends, we consider only simulators that support the same number of qubits as the real backends. Consequently, we selected simulators that support 127 qubits, namely FakeWashington and FakeSherbrooke. In summary, we consider the following backends:
4.2.1 Quantum Computers
@@ 4.2.2 Quantum Simulators
4.3 Measurement of execution time
In order to understand the execution time of quantum programs, build and evaluate our approach for quantum execution time estimation, we execute our subject quantum programs (see 4.1) on the selected backends (see 4.2) and measure the execution time.
Measurement of execution time on simulators. When users want to run a quantum program on a simulator, they need to select a simulator (e.g., FakeSherbrooke) and compile the program with the simulator as the compilation target. Then, the compiled program should be submitted to the simulator for execution. The total time consumed in the process of execution is what we measure as the execution time of a quantum program on a simulator. Qiskit provides an API [^3] https://docs.quantum.ibm.com/api/qiskit/0.45/execute that makes it easy to obtain the time used by a simulator to execute a quantum program. Each quantum program is executed for 1024 shots on each simulator.
Measurement of execution time on quantum computers. Similarly, when users choose to execute a quantum program on a quantum computer, such as using the ibm_osaka provided by the IBM Quantum Platform, they need to submit the designed and compiled quantum program to the IBM Quantum Platform, select ibm_osaka as the quantum computer, and then queue for its execution. It is important to note that in order to select quantum computers provided by the IBM Quantum Platform to execute a quantum program, a token must be provided. Every IBM Quantum Platform account possesses this token, which grants access to the quantum computers on the platform. This token allows users to submit the designed and compiled quantum program to a specified quantum computer for execution.
Finally, the total time spent during the execution process is what we measure as the execution time of a quantum program on a quantum computer. The process of retrieving the execution time is similar to that on a simulator. We use the same API to obtain the execution time of quantum circuits on quantum computers as we do for simulators.
To ensure consistency in results, each quantum program is also executed for 1024 shots on quantum computers.
4.4 Determining minimum repeated measures for quantum execution time accuracy
In order to compensate for measurement errors, we ascertain the lowest requisite number of repeated measures for the execution time of each quantum circuit to maintain a satisfactory precision level, as delineated by Equation  1 [8].
$n=\left\lceil\left(\frac{100\times z_{1-\alpha/2}\times s}{r\times\bar{x}}\right)^{2}\right\rceil$ (1)
@@ 4.4 Determining minimum repeated measures for quantum execution time accuracy
5 Quantum execution time prediction model
In this section, we describe how we build and evaluate our quantum execution time prediction model, including the representation of the quantum circuits, the construction of the prediction model, the active learning for collecting training data for quantum programs running on real quantum computers, as well our evaluation method and metrics.
5.1 Quantum circuit representation
@@
Additionally, based on previous related research, we also consider five other features to characterize the global attributes of each quantum circuit. These features include program communication [3], critical-depth, entanglement-ratio [6], parallelism [14], and liveness [32], as originally introduced in prior work [28]. This methodical approach allows for a nuanced and detailed representation of the quantum circuits, facilitating more accurate and efficient training of the machine learning model.
Program Communication measures the amount of interaction required between qubits in a quantum circuit. It is quantified by calculating the average degree of interaction between qubits, where a qubit’s ”degree” refers to the number of other qubits it interacts with through multi-qubit operations [11]. This measure helps assess how communication-intensive a quantum program is: sparsely connected applications will have values near zero, while denser programs will be close to one.
Critical-Path in a quantum circuit is the longest chain of dependent operations from start to finish. This path is crucial because it involves two-qubit operations, which are more error-prone and time-consuming [16]. The critical depth, calculated as the ratio of two-qubit operations on this path to the total in the circuit, indicates how sequentially dependent the circuit is; a value close to 1 means the circuit is highly serialized.
@@
Parallelism refers to the degree to which quantum operations can be performed simultaneously. Parallel operations can speed up execution [14]. The parallelism feature is determined by comparing the ratios of the number of qubits, the number of gates, and the circuit depth. This comparison helps assess how much parallelism is present in the circuit. Highly parallel circuits, which perform many operations in parallel, will have a parallelism feature value close to 1.
Liveness in a quantum circuit refers to the state of qubits during program execution. A qubit can either be actively involved in computation or remain idle, waiting for the next instruction. Ideally, a qubit should maintain its state while idling, but in reality, environmental interactions can cause the qubit to lose coherence  [33]. The liveness feature captures how often qubits are actively used versus how often they are idle during the execution of a quantum circuit. This feature is calculated by forming a liveness matrix. In this matrix, each row represents a qubit, and each column corresponds to a time step in the circuit’s execution. The matrix entries are 1 if the qubit is active and 0 if it is idle.
@@
Figure 3: An example of the global features of a quantum circuit (Dimensionality: 41)
5.1.2 graph features
A quantum circuit can be conceptualized as a graph, allowing for the extraction of graph-based information for analytical purposes. The Graph Transformer [31, 37], derived from attention-based [30] Transformer models renowned for their effectiveness in sequence modeling, represents a robust method for this type of analysis. By leveraging the attention mechanism, the Graph Transformer updates the features of each node in successive layers. This approach has been applied in recent studies to estimate the fidelity of quantum circuits, where each node in the graph represents a qubit, quantum gate, or measurement, and edges denote the time-dependent sequence of operations [34]. This graph-based representation facilitates a nuanced understanding of the circuit’s dynamics and interactions.
@@
Figure 4: An example of converting a quantum circuit into a directed acyclic graph (dag)
@@
In our experimentation, we align the maximum number of qubits to 127, mirroring the upper limit supported by contemporary quantum hardware, resulting in feature vectors spanning 178 elements. The leading 46 entries constitute a one-hot encoding for the node classification, including the initial qubit state, measurements, and 44 distinct gate types. Subsequently, 127 entries demarcate the qubit(s)’ positions, with a ”1” denoting the qubit’s involvement in a gate operation. This scheme extends to gates influencing multiple qubits. Calibration details for the backend are represented by four entries formatted as [T1, T2 for the primary target qubit, T1, T2 for the secondary target qubit]. Inapplicable features for specific nodes are assigned a zero value. For instance, a gate like X, which impacts only one qubit, would have the T1 and T2 values for a second qubit set to zero. The final entry of the vector is reserved for the node’s index. Fig. 5 illustrates the complete composition of the feature vector.
Figure 5: An example node of the graph features of a quantum circuit (Dimensionality: 178)
@@
5.2 Model construction for execution time prediction

### v2

@@ 2.3. Classical software execution time estimation
3. Background
3.1. Quantum circuits
Fig. 1 shows a simple quantum circuit that we developed using IBM’s open-source software development toolkit, Qiskit. A quantum circuit typically consists of a combination of quantum bits (qubits), classical bits, quantum gates, and measurement gates. In the example shown in Fig. 1, $q_{0}$ and $q_{1}$ represent two qubits, whereas $c$ represent two classical bits. Unlike classical bits, each qubit can exist simultaneously in both $|0\rangle$ and $|1\rangle$ states.
@@ 3.1. Quantum circuits
  • The blue box with ”H” symbol: Hadamard gate, applied to $q_{0}$.
  • The blue box with ”+” symbol: CNOT gate, with $q_{0}$ as the control qubit and $q_{1}$ as the target qubit.
@@
Figure 1. An example of a Quantum Circuit
3.2. Quantum circuit compilation
@@
Like the classical domain, implementing a quantum algorithm on a real backend involves transforming it into a format that meets the specific hardware constraints (quetschlich2023mqt). This transformation encompasses four stages: the algorithmic level, target-independent level, target-dependent native gates level, and target-dependent mapped level.
@@
Quantum algorithms are typically designed and trialed at a hardware-neutral stage, known as the algorithmic level (quetschlich2023mqt). Here, they are expressed as quantum circuits made up of high-level elements, free of restrictions related to specific gate-sets or hardware designs.
@@
The initial step in applying a conceptual quantum algorithm to a specific problem is to create and refine these high-level components, disregarding the intended hardware. This process is akin to the function of a classical compiler and includes optimizing and simplifying the representation, now termed the target-independent level (quetschlich2023mqt).
@@
Modern quantum backends have specific limitations, necessitating a target-dependent compilation phase. Backends offer a limited range of native gates, usually an entangling gate and a set of single-qubit gates. Circuits must therefore be adapted to this native gate-set and optimized to minimize additional complexity. This adaptation is known as the target-dependent native gates level (quetschlich2023mqt).
@@
Additionally, current backends, particularly those using superconducting qubits, have restricted qubit connectivity. This requires mapping the circuit’s logical qubits to the backend’s physical qubits, ensuring that multi-qubit gates apply only to directly connected qubits. This dynamic mapping, followed by optimization to reduce compilation overhead, is defined as the target-dependent mapped level (quetschlich2023mqt).
@@
Finally, the circuit is prepared for execution on the quantum computer. It involves scheduling the gates, calibrating, and converting the circuit into machine code (quetschlich2023mqt). The final binary code is then executed on the backend to perform the quantum computation.
3.3. Quantum execution and execution time
@@
When a quantum circuit is submitted for execution on a quantum computing platform such as the IBM Quantum Platform, the platform usually provides an estimated execution time. The execution time includes loading the quantum circuit into the pre-selected quantum computer that creates the pulses, running multiple repetitions of calibration pulses and the quantum circuit, and resetting the qubits (relaxation) along with either calibration or executing the quantum circuit. These steps together make up the total execution time. As illustrated in Fig. 2(a), on the IBM quantum computer ibm_osaka, the quantum circuit is expected to run for 2 minutes and 27.5 seconds. After the execution, the platform provides the actual execution time of the circuit. As shown in Fig. 2(b), the actual execution time is only 27 seconds, substantially different from the estimation. The primary objective of our approach is to accurately predict the execution time, aiming to substantially reduce the significant errors observed in the estimates provided by quantum computing platforms. Please note that currently, IBM Quantum Platform only provides the worst case estimated execution time, which serves primarily as an upper bound for latency-critical scenarios. However, this estimate alone is insufficient for many practical use cases where developers and platform operators require more realistic predictions. Our method is expected to complement IBM’s worst-case latency estimates by providing precise average-case execution times, which are essential for realistic cost estimation; developers can rely on IBM’s upper bounds when strict real-time guarantees are required and leverage our predictions for practical runtime planning. While the IBM Quantum Platform serves as a case study in this paper, our goal extends beyond this to a more general application across various platforms.
@@
Figure 2. Estimated and Actual Execution Times on IBM Quantum Platform
4. Data Preparation
In this section, we describe our data preparation process, including the selection of the quantum circuits and quantum backends, as well as the measurement of execution time.
4.1. Quantum circuit selection
We select our subject quantum circuits from the MQT Benchmark (Quetschlich2023mqtbench) which provides an extensive collection of quantum circuits covering various quantum computing applications across different levels of abstraction. This library comprises 3,000 quantum circuits, spanning from 2 to 130 qubits. Notably, this dataset exhibits significant diversity as it considers two qubit technologies, seven different backends, two compilers, and six corresponding settings. This diversity renders it a robust and versatile resource for our research that aims to estimate quantum execution time for diverse quantum circuits. Due to the limitation of quantum computing resources, we must select a subset for our experiments. For the compiler, we have chosen Qiskit, an open-source SDK tailored for quantum computing. Qiskit specializes in circuit, pulse, and algorithm levels, enabling the creation and manipulation of quantum circuits. These circuits can be executed on prototype quantum backends available through the IBM Quantum Platform or via simulators on local machines. Adhering to the circuit model of universal quantum computation, Qiskit is versatile, supporting various quantum hardware types that comply with this model, including superconducting qubits and trapped ions.
@@ 4.1. Quantum circuit selection
Finally, as for the quantum circuits, we selected all quantum circuits ranging from 2 qubits to 127 qubits (the limit of our considered quantum backends). When executing these quantum circuits, we found that a small portion encountered loading or compilation errors, and some had execution times that were excessively long, exceeding 10 minutes. Therefore, we excluded these quantum circuits with errors and set a timeout function to filter out those with excessively long execution times (over 10 minutes). Ultimately, we obtained a total of 1,510 subject quantum circuits.
4.2. Quantum backend selection
We consider the quantum backends provided by the IBM Quantum Platform as it is one of the mostly used quantum computing platforms and provides open accesses for free. Importantly, all the real devices and corresponding simulators used in this study are based on the superconducting qubit technology, while other hardware modalities (e.g., ion-trap processors) are outside the scope of our evaluation. For open accesses, the accessible quantum computers include ibm_brisbane, ibm_osaka, and ibm_kyoto (at the time of writing), all of which support 127 physical qubits. Due to the high traffic on ibm_brisbane and in order to ensure data diversity, we have chosen to execute the quantum circuits on both ibm_osaka and ibm_kyoto. To maintain consistency in data dimensions while leveraging simulation data and active learning to select samples for execution on real backends, we consider only simulators that support the same number of qubits as the real backends. Consequently, we selected simulators that support 127 qubits, namely FakeWashington and FakeSherbrooke. In summary, we consider the following backends:
4.2.1. Quantum Computers
@@ 4.2.2. Quantum Simulators
4.3. Measurement of execution time
In order to understand the execution time of quantum circuits, build and evaluate our approach for quantum execution time estimation, we execute our subject quantum circuits (see 4.1) on the selected backends (see 4.2) and measure the execution time.
Measurement of execution time on simulators. When users want to run a quantum circuit on a simulator, they need to select a simulator (e.g., FakeSherbrooke) and compile the circuit with the simulator as the compilation target. Then, the compiled circuit should be submitted to the simulator for execution. The total time consumed in the process of execution is what we measure as the execution time of a quantum circuit on a simulator. Qiskit provides an API [^4] https://docs.quantum.ibm.com/api/qiskit/0.45/execute that makes it easy to obtain the time used by a simulator to execute a quantum circuit. Each quantum circuit is executed for 1024 shots on each simulator.
Measurement of execution time on quantum computers. Similarly, when users choose to execute a quantum circuit on a quantum computer, such as using the ibm_osaka provided by the IBM Quantum Platform, they need to submit the designed and compiled quantum circuit to the IBM Quantum Platform, select ibm_osaka as the quantum computer, and then queue for its execution. It is important to note that in order to select quantum computers provided by the IBM Quantum Platform to execute a quantum circuit, a token must be provided. Every IBM Quantum Platform account possesses this token, which grants access to the quantum computers on the platform. This token allows users to submit the designed and compiled quantum circuit to a specified quantum computer for execution.
Finally, the total time spent during the execution process is what we measure as the execution time of a quantum circuit on a quantum computer. The process of retrieving the execution time is similar to that on a simulator. We use the same API to obtain the execution time of quantum circuits on quantum computers as we do for simulators.
To ensure consistency in results, each quantum circuit is also executed for 1024 shots on quantum computers.
4.4. Determining minimum repeated measures for quantum execution time accuracy
In order to compensate for measurement errors, we ascertain the lowest requisite number of repeated measures for the execution time of each quantum circuit to maintain a satisfactory precision level, as delineated by Equation  1 (bookJainRaj).
(1) $n=\left\lceil\left(\frac{100\times z_{1-\alpha/2}\times s}{r\times\bar{x}}\right)^{2}\right\rceil$
@@ 4.4. Determining minimum repeated measures for quantum execution time accuracy
5. Quantum execution time prediction model
In this section, we describe how we build and evaluate our quantum execution time prediction model, including the representation of the quantum circuits, the construction of the prediction model, the active learning for collecting training data for quantum circuits running on real quantum computers, as well our evaluation method and metrics.
5.1. Quantum circuit representation
@@ 5.1.1. Global features
Additionally, based on previous related research, we also consider five other features to characterize the global attributes of each quantum circuit. These features include program communication (black2002quantum), critical-depth, entanglement-ratio (GUZMAN2004382), parallelism (murali2021instruction), and liveness (viola1999dynamical), as originally introduced in prior work (tomesh2022supermarq). This methodical approach allows for a nuanced and detailed representation of the quantum circuits, facilitating more accurate and efficient training of the machine learning model.
Program Communication measures the amount of interaction required between qubits in a quantum circuit. It is quantified by calculating the average degree of interaction between qubits, where a qubit’s ”degree” refers to the number of other qubits it interacts with through multi-qubit operations (li2019tackling). This measure helps assess how communication-intensive a quantum program is: sparsely connected applications will have values near zero, while denser programs will be close to one.
Critical-Path in a quantum circuit is the longest chain of dependent operations from start to finish. This path is crucial because it involves two-qubit operations, which are more error-prone and time-consuming (erata2024quantum; 9678792). The critical depth, calculated as the ratio of two-qubit operations on this path to the total in the circuit, indicates how sequentially dependent the circuit is; a value close to 1 means the circuit is highly serialized.
@@
Parallelism refers to the degree to which quantum operations can be performed simultaneously. Parallel operations can speed up execution (murali2021instruction). The parallelism feature is determined by comparing the ratios of the number of qubits ($n$), the number of gates ($n_{g}$), and the circuit depth ($d$):
(2) $P=\left(\frac{n_{g}}{d}-1\right)\cdot\frac{1}{n-1}$
This comparison helps assess how much parallelism is present in the circuit. Highly parallel circuits, which perform many operations in parallel, will have a parallelism feature value close to 1.
Liveness in a quantum circuit refers to the state of qubits during program execution. A qubit can either be actively involved in computation or remain idle, waiting for the next instruction. Ideally, a qubit should maintain its state while idling, but in reality, environmental interactions can cause the qubit to lose coherence  (viola1999dynamical). The liveness feature captures how often qubits are actively used versus how often they are idle during the execution of a quantum circuit. This feature is calculated by forming a liveness matrix. In this matrix, each row represents a qubit, and each column corresponds to a time step in the circuit’s execution. The matrix entries are 1 if the qubit is active and 0 if it is idle. To convert this matrix into a single scalar value representing liveness, we compute the average number of active qubit-time pairs. This is done by summing all entries in the liveness matrix and dividing by the total number of elements, which is the product of the number of qubits and the circuit depth. The value ranges from 0 to 1, where a higher value indicates more active usage of qubits during execution.
@@
Figure 3. An example of the global features of a quantum circuit (Dimensionality: 41)
5.1.2. Graph features
A quantum circuit can be conceptualized as a graph, allowing for the extraction of graph-based information for analytical purposes. The Graph Transformer (velickovic2017graph; yun2019graph), derived from attention-based (vaswani2017attention) Transformer models renowned for their effectiveness in sequence modeling, represents a robust method for this type of analysis. By leveraging the attention mechanism, the Graph Transformer updates the features of each node in successive layers. This approach has been applied in recent studies to estimate the fidelity of quantum circuits, where each node in the graph represents a qubit, quantum gate, or measurement, and edges denote the time-dependent sequence of operations (wang2022quest). This graph-based representation facilitates a nuanced understanding of the circuit’s dynamics and interactions.
@@ 5.1.2. Graph features
Figure 4. An example of converting a quantum circuit into a directed acyclic graph (dag)
@@
In our experimentation, we align the maximum number of qubits to 127, mirroring the upper limit supported by contemporary quantum hardware, resulting in feature vectors spanning 178 elements. The leading 46 entries constitute a one-hot encoding for the node classification, including the initial qubit state, measurements, and 44 distinct gate types. Subsequently, 127 entries demarcate the qubit(s)’ positions, with a ”1” denoting the qubit’s involvement in a gate operation. This scheme extends to gates influencing multiple qubits. Note that the quantum circuits used in this experiment include only single- and two-qubit gates. Calibration details for the backend are represented by four entries formatted as [T1, T2 for the primary target qubit, T1, T2 for the secondary target qubit]. Inapplicable features for specific nodes are assigned a zero value. For instance, a gate like X, which impacts only one qubit, would have the T1 and T2 values for a second qubit set to zero. The final entry of the vector is reserved for the node’s index. Fig. 5 illustrates the complete composition of the feature vector.
Figure 5. An example node of the graph features of a quantum circuit (Dimensionality: 178)
@@
5.2. Model construction for execution time prediction